OpenStudy (anonymous):

Two lines form a linear system. The slope of one line is 4 and it has a y-intercept of (0, -3). The slope of the other line is 4 and it has a y-intercept of (0, 8). How many solutions does this system have?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

your a ugly feather. bye now.

OpenStudy (saifoo.khan):

Lol. Jokes apart.

OpenStudy (bahrom7893):

Saifoo, no need to insult and then delete the responses.

OpenStudy (saifoo.khan):

Line 1: y = mx + b y = 4x - 3 Line 2: y = mx + b y = 4x + 8

OpenStudy (anonymous):

i know right, hes rude. Thank youuu

OpenStudy (saifoo.khan):

Where "m" is the Slope and b is the y intercept.

OpenStudy (saifoo.khan):

@bahrom, I only deleted the Abusive words that's it.

OpenStudy (bahrom7893):

Yea, well you shouldn't be posting it in the first place.. and hey abby if he called u dumb, don't worry about it. He can only answer up to like 9th grade problems anyway.

OpenStudy (saifoo.khan):

Lmao, i never called her that. I Dont call people Dumb. :p

OpenStudy (anonymous):

well you posted a rude video. SO GO SOMEWHERE PLEASE.

OpenStudy (anonymous):

haha wow. & thanks bahrom (:

OpenStudy (anonymous):

But can someone answer my question tho? (:

OpenStudy (bahrom7893):

what's the question?

OpenStudy (saifoo.khan):

Bahrom, Ultimate Enemy!

OpenStudy (saifoo.khan):

I just answered it though @bieber

OpenStudy (bahrom7893):

lol don't worry mate, u're a level below haha.

OpenStudy (saifoo.khan):

Hahah, Good Point^

OpenStudy (anonymous):

Two lines form a linear system. The slope of one line is 4 and it has a y-intercept of (0, -3). The slope of the other line is 4 and it has a y-intercept of (0, 8). How many solutions does this system have?

OpenStudy (saifoo.khan):

Line 1: y = mx + b y = 4x - 3 Line 2: y = mx + b y = 4x + 8 Now solve both line 1 and line 2.

OpenStudy (anonymous):

i dont want your answer saifoo, u sent me that video..? go away.

OpenStudy (saifoo.khan):

Lol, Sure. As you wish.

OpenStudy (saifoo.khan):

Need help now? Or i should stay away?

OpenStudy (bahrom7893):

hey no need to abuse the fact that u know something that she doesn't

OpenStudy (saifoo.khan):

She dont want to learn.. now what can is say..

OpenStudy (bahrom7893):

abbey post in new threads, nobody comes back to old posts

OpenStudy (anonymous):

someone just tell me the answer pls

OpenStudy (saifoo.khan):

Someone (not me) She's talking to you Bahrom.

OpenStudy (bahrom7893):

the answer to what? saifoo already answered it.. ok lol

OpenStudy (bahrom7893):

Two lines form a linear system. The slope of one line is 4 and it has a y-intercept of (0, -3). The slope of the other line is 4 and it has a y-intercept of (0, 8). How many solutions does this system have? y = 4x - 3 y = 4x + 8 4x-3 = 4x + 8 -3 != 8 So no solutions

OpenStudy (anonymous):

okay! Thaaank you (: